# Exploring Kolmogorov-Arnold Networks for Interpretable Time Series Classification

### v2

@@ III-B2 MLP FLOPs computation
For a neuron with $d_{\text{in}}$ inputs, calculating the output requires $2\times d_{\text{in}}$ FLOPs (for weights) and 1 FLOP for the bias, resulting in a total of $12\times d_{\text{in}}+$ 1 FLOPs. If there are $d_{\text{out}}$ output neurons, the total FLOPs for the layer is $d_{\text{out}}\times(2\times d_{\text{in}}+1)$.
We consider fully connected MLP which has input layer with $T$ neurons, $K$ hidden layers, with $M$ neurons in each, and $C$ neurons in output layer. FLOPs for forward propagation are therefore calculated as
$N_{FP}=(M+2MT)+(K-1)(M+2M^{2})+(M+2MC)$ (11)
@@ IV Methodology
In this section we elaborate on the methodology adopted for this study ($https://github.com/irina-b1/KAN_{T}S$). First we provide considerations on data and preprocessing, followed by reference model configurations, training process, hyperparamater, complexity and interpretability analysis.
@@ IV-A Data and preprocessing
In this study, the UCR (University of California, Riverside)[29] dataset archive time series classification benchmark is utilized. The datasets were selected based on their completeness, since some datasets include time series instances containing missing values. To ensure that an adequate amount of data remained for both the training and testing phases, these datasets were excluded, resulting in total of 117 datasets used.
In terms of size, the datasets vary significantly, with some containing 16 time series instances, while others encompassed up to 8,926 instances. The input lengths of these time series also showed considerable diversity, ranging from 15 to 2,844 time steps per instance in different datasets. Furthermore, the datasets included varying numbers of classes for classification tasks, with scenarios ranging from binary classifications with 2 classes to multiclass problems with up to 60 distinct classes.
@@ IV-B Training setup
For training, each dataset was partitioned into into a training dataset and a validation dataset in an 80:20 ratio. The test datasets are provided separately by default.
The training process across all models used the Adam optimizer with a learning rate of $0.001$. However, instead of the mean absolute error (MAE) loss function used for regression tasks in the original work, we employed the cross-entropy loss function to address the classification nature of our problem. Additionally, we introduced $L1$ regularization with a weight factor of $0.1$.
Similarly, in the MLP models, we used the same Adam optimizer, learning rate ($0.001$), and cross-entropy loss function as in the KAN models, for the same reasons. However, we applied L2 regularization with a weight factor of 1, selected through empirical tuning, instead of L1 regularization.
To ensure robustness and mitigate the effects of random initialization, each model was trained on 5 different random seeds. The results presented are averages of these 5 runs. The models were trained for 500 epochs using batch size 16.
The original KAN library [14] was modified to enable GPU support, allowing the experiments to leverage the computational power of GPUs rather than relying solely on CPU execution. Thus, all evaluations involving KAN models were conducted on an NVIDIA A100 80GB GPU.
For performance evaluation, we use precision, recall and F1 score. The precision measures how many instances predicted as a certain class actually belong to that class, expressed as: $\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$, where TP (true positives) represents the number of correctly identified positive instances, and FP (false positives) refers to the number of instances incorrectly identified as positive. Recall measures how many instances of some class were correctly detected, given by: $\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$, where FN (false negatives) represents the number of actual positive instances that the model failed to identify as positive. The F1 score is expressed as: $F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$, where higher values indicate better performance in balancing precision and recall (better classification).
@@ IV-C Reference architecture configurations
Initially, we analyse the performance of KAN and MLP architectures originally proposed for forecasting by [23]. The goal is to gain insights into the generalization capabilities and transfer potential of these architectures from regression to classification problems.
As discussed in Section III, we employ two different models: KAN and MLP. Alongside the official implementation of KAN 111github.com/KindXiaoming/pykan, we also include its variant, Efficient KAN 222github.com/Blealtan/efficient-kan, which optimizes memory by reformulating the activation process, applying B-splines directly to inputs and combining them linearly, avoiding the need for large tensor expansions. It replaces input-based L1 regularization with weight-based L1, improving efficiency. The L1 regularization is now computed as mean absolute value of the spline weights [30].
KAN and Efficient KAN are configured as [$T$, 40, 40, $C$], representing a 3-depth network, where $T$ represents the length of the input time series, while $C$ denotes number of classes for classification, as per Section III. A 4-depth variant includes an additional hidden layer with 40 nodes. We retained the spline order $k=3$ and a fixed grid size $G=5$.
Similarly, the MLP architecture is configured as [$T$, 300, 300, 300, $C$], representing a 3-depth network. The 4-depth variant adds a hidden layer of $300$ nodes.
@@ IV-D Hyperparameter impact analysis
To provide deeper insights into the KAN architecture’s performance, we conduct an analysis of hyperparameters and configuration parameters with the goal of finding the best model for time series classification on the UCR behcmark. We conducted a series of hyperparameter impact analyses by varying key hyperparameters, such as grid size $G$, depth (number of layers) $L$, and layer size (number of nodes per layer) $M$. For each of these evaluations, the models were trained using different learning rates $lr\in\{0.0001,0.001,0.01,0.1,1\}$.
@@ IV-E Complexity analysis
We evaluated resource consumption by calculating the average FLOPs over 117 datasets for one prediction, and the theoretical energy consumption (TEC) per prediction. FLOPs were computed theoretically using Eq. 8 and 11 outlined in Section III. As per Section IV-A, $T\in\{16,\ldots,8,926\}$, $C\in\{2,\ldots,60\}$.
We estimated TEC using the Eq. $\text{TEC}=\frac{\text{FLOPs}}{\text{FLOPS/Watt}}$ [31], where FLOPS/Watt represents the number of floating-point operations executed per second per watt. As the experiments were conducted on an NVIDIA A100 80GB PCIe GPU, its theoretical FLOPS/Wat of 65 GFLOPS, for float32 operations, was used for calculation.
@@ IV-F Interpretability analysis
Furthermore, to gain further insight into the decision making of KAN, global SHapley Additive exPlanations (SHAP) values [32] were used. SHAP offers a model-agnostic framework for global interpretability by quantifying the average contribution of each feature to the model’s predictions across the entire dataset. The contribution of a feature $x_{i}$ is represented by its Shapley value, $\phi_{i}$, which is computed based on cooperative game theory. For a model the prediction for an instance $X$ is expressed as:
$SHAP=\phi_{0}+\sum_{i=1}^{M}\phi_{i}$ (12)
where $\phi_{0}$ is the model’s baseline expected value across the dataset, and $\phi_{i}$ is the Shapley value of feature $x_{i}$, and M represents the total number of features. By aggregating these values SHAP enables insights into the global decisions of the model. In addition SHAP is also utilized to compare the feature importance of KANs to the MLP architecture.
@@
TABLE I: Performance of reference architectures introduced in [23] for forecasting.
@@ V-A Classification performance analysis of reference architectures
Table I presents the results of the classifiers. The table is divided into rows representing different models and their configurations, with the first two rows listing the results for MLP models and the subsequent four rows listing the results for KAN and Efficient KAN models. The first column lists the type of model, while the second column details the specific configuration of each model, including the nodes per layer and grid size for KAN models. The following columns provide the performance metrics for each model, split into three main categories: precision, recall, and F1 score. Each of these categories is further divided into two subcolumns that show the mean and standard deviation (StdDev) of the results across multiple runs. The final column presents average time, measured in seconds, for training one model.
Looking at the first two rows of the table, the MLP models demonstrate relatively high and consistent performance across precision, recall, and F1 score metrics in both depths. For the 3-depth MLP, the mean precision is 0.77 with a standard deviation of 0.18, while the 4-depth model exhibits a similar precision mean of 0.76 and a slightly lower standard deviation of 0.16. In terms of recall, both depths show identical mean values of 0.64 and a standard deviation of 0.26, indicating no change in the model’s ability to identify relevant instances with increased depth. Additionally, both models maintain the same F1 score mean of 0.62, although the standard deviation slightly increases from 0.26 in the 3-depth model to 0.27 in the 4-depth model, suggesting a minor decrease in consistency for the F1 score with the deeper model.
Rows 3 and 4 for the original KAN implementation show a substantial drop in performance compared to the MLP models. The 3-depth KAN has a mean precision of 0.38, with a recall of 0.33 and an F1 score of 0.30. The 4-depth KAN model continues this trend with slightly lower performance, showing a precision of 0.37, recall of 0.32, and F1 score of 0.29. The further decline in performance suggests that adding depth does not improve generalization but may instead lead to overfitting, especially across multiple diverse datasets. These results indicate that KAN struggles to match the classification performance of the MLP. While previous research demonstrated that KAN performed well for time series prediction and exhibited strong generalization abilities [23], here KAN’s transferability to a different problem, classification, appears less successful.
Lower performance results come with a significantly reduced training time of $\sim 45$ seconds for 3-depth, due to the GPU acceleration, as described in Section IV-E, suggesting a trade-off between performance and efficiency. However, the training time in 4-depth increases to $\sim 57$ seconds, reflecting the impact of the added depth on computational cost.
@@
TABLE II: Comparison of performance and computational characteristics across the reference models
The last two rows show results for the models trained with the Efficient KAN implementation, which offer a notable improvement in performance compared to the original KAN implementation. The Efficient KAN with 3-depth, demonstrates a significant boost in performance with a precision of 0.72, recall of 0.70, and F1 score of 0.70. Similarly, the 4-depth Efficient KAN, maintains high performance metrics with a precision of 0.71, recall of 0.70, and F1 score of 0.69. Improvements in comparison to KAN can be attributed to implementation changes, as discussed in Section IV-C.
@@ V-B Complexity analysis for the reference models
Table II presents the computational complexity analysis for the MLP and KAN configurations. The first two columns, listing the model and configuration, are referenced from Table I. The subsequent columns provide key complexity metrics: the third column displays the total number of lernable parameters in each model, while the fourth column reports the average theoretical FLOPs for one prediction, based on model architecture discussed in Section III, across all used UCR datasets. The last column presents the Theoretical Energy Consumption (TEC) in Joules, calculated based on the FLOPs and GPU efficiency, as per Section IV-E.
Examining the first two rows of the table, we observe that MLPs, with approximately $344k$ and $434k$ learnable parameters for depths $3$ and $4$ respectively, are the most computationally demanding models in terms of learnable parameters and training duration. Their theoretical FLOPs reach $688,420$ and $868,720$, while requiring $1.008\times 10^{-5}$ and $1.336\times 10^{-5}$ Joules of energy per single prediction, positioning them as the best-performing models in theoretical FLOPs and TEC (Total Energy Consumption).
Rows 3-4 of the table reveal that the original KAN implementation has noticeably fewer learnable parameters compared to the MLP models, consistent with the theoretical analysis in Section III-A. Despite the reduction in parameters, they come with significantly higher theoretical FLOPs, with values of $6,146,993$ for depth 3 and $6,566,993$ for depth 4, due to B-spline computation, as per Eq. 8. The energy required for a single prediction also reflects this increase, amounting to $9.457\times 10^{-5}$ Joules for depth 3 and $10.103\times 10^{-5}$ Joules for depth 4. These results illustrate a trade-off, where KAN achieves a smaller parameter count but requires greater computational resources in terms of FLOPs and energy per prediction.
The last two rows for the Efficient KAN implementation show the same number of weights and theoretical FLOPs as the original implementation, as the model architectures are identical. Despite higher energy consumption, Efficient KAN maintains fast training times. Along with the discussed increased performance, this balance of high performance and lower computational costs makes these models strong candidates for applications requiring both efficiency and effectiveness.

### v3

@@ III-B2 MLP FLOPs computation
For a neuron with $d_{\text{in}}$ inputs, calculating the output requires $2\times d_{\text{in}}$ FLOPs (for weights) and 1 FLOP for the bias, resulting in a total of $2\times d_{\text{in}}+1$ FLOPs. If there are $d_{\text{out}}$ output neurons, the total FLOPs for the layer is $d_{\text{out}}\times(2\times d_{\text{in}}+1)$.
We consider fully connected MLP which has input layer with $T$ neurons, $K$ hidden layers, with $M$ neurons in each, and $C$ neurons in output layer. FLOPs for forward propagation are therefore calculated as:
$N_{FP}=(M+2MT)+(K-1)(M+2M^{2})+(M+2MC).$ (10)
@@ IV Methodology
In this section we elaborate on the methodology adopted for this study [^1] https://github.com/irina-b1/KAN_TS. First we provide considerations on data and preprocessing, followed by reference model configurations, training process, hyperparamater, complexity and interpretability analysis.
@@ IV-A Data and preprocessing
In this study, the UCR (University of California, Riverside) [28] dataset archive time series classification benchmark is utilized , which contains a total of 128 univariate datasets from diverse domains, ranging from ECG signals and motion capture data to spectrographs and simulated control systems.
The datasets were selected based on their completeness, since some datasets include time series instances containing missing values. To ensure that an adequate amount of data remained for both the training and testing phases, these datasets were excluded, resulting in total of 117 datasets used. For each dataset, the Appendix provides the number of training and test instances, the length of time series, and the number of classes.
Table I further shows descriptive statistics for 117 datasets. Each row represents a statistical measure: minimum, maximum, mean, median and standard deviation. The columns correspond to the number of training samples, the number of test samples, the length of each time series and the number of classes across the 117 UCR datasets.
In terms of size, for training specifically, the datasets vary significantly, ranging from 16 up to 8,926 time series instances. The input lengths of these time series also showed considerable diversity, ranging from 15 to 2,844 time steps per instance in different datasets. Furthermore, the datasets included varying numbers of classes for classification tasks, with scenarios ranging from binary classifications with 2 classes to multiclass problems with up to 60 distinct classes. To illustrate the diversity in shape and class structure, we visualize 15 randomly selected datasets in Fig. 3 , where each subplot displays one representative instance per class, with distinct colors denoting different classes.
@@
Data Availability Statement. The data that support the findings of this study are available in the UCR Archive at https://doi.org/10.1109/JAS.2019.1911747, and are also available in the public domain at https://www.timeseriesclassification.com/.
Figure 3: 15 datasets showing one representative time series per class. Distinct colors indicate different classes
Statistic Train Test Length Class
Min 16 20 15 2
Max 8 926 16 800 2 844 60
Mean 496.72 1 086.72 537.10 8.26
Median 181 343 301 3
StdDev 1 155.16 2 080.59 583.14 12.26
TABLE I: Summary statistics of the 117 UCR archive time series datasets
@@ IV-B Training setup
For training, each dataset was partitioned into a training dataset and a validation dataset in an 80:20 ratio. The test datasets are provided separately by default. To ensure robustness and mitigate the effects of random initialization, each model was trained on 5 different random seeds. The presented results are averages of these 5 runs across all 117 datasets. The models were trained for 500 epochs using batch size 16.
The training process across all models used the Adam optimizer. The baseline architectures from [23] were trained with a learning rate of $0.001$. However, instead of the mean absolute error (MAE) loss function used for regression tasks in the original work, we employed the cross-entropy loss function to address the classification nature of our problem. Additionally, we introduced $L1$ regularization with a weight factor of $0.1$.
Similarly, in the baseline MLP models, we used the same Adam optimizer, learning rate ($0.001$), and cross-entropy loss function as in the KAN models, for the same reasons. However, instead of L1 we applied L2 regularization with a weight factor of 1, selected through empirical tuning.
The original KAN library [14] was modified to enable GPU support, allowing the experiments to leverage the computational power of GPUs rather than relying solely on CPU execution. Thus, all evaluations of each model were conducted on an NVIDIA A100 80GB GPU.
For performance evaluation, we use precision, recall and F1 score. The precision measures how many instances predicted as a certain class actually belong to that class, expressed as: $\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$, where TP (true positives) represents the number of correctly identified positive instances, and FP (false positives) refers to the number of instances incorrectly identified as positive. Recall measures how many instances of some class were correctly detected. It is given by: $\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$, where FN (false negatives) represents the number of actual positive instances that the model failed to identify as positive. The F1 score is expressed as: $F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$, where higher values indicate better performance in balancing precision and recall (better classification).
@@ IV-C Reference architecture configurations
Initially, we analyze the performance of KAN and MLP by adopting the regression designs from the extensive empirical study in [23]. The goal is to gain insights into the generalization capabilities and transfer potential of these architectures from regression to classification problems.
As discussed in Section III, we employ two different models: KAN and MLP. Alongside the original implementation of KAN222github.com/KindXiaoming/pykan, we also include its variant, Efficient KAN333github.com/Blealtan/efficient-kan, which optimizes memory by reformulating the activation process, applying B-splines directly to inputs and combining them linearly, thus avoiding the need for large tensor expansions. It replaces input-based L1 regularization with weight-based L1, improving efficiency. The L1 regularization is now computed as mean absolute value of the spline weights [29].
KAN and Efficient KAN are configured as [$T$, 40, 40, $C$], representing a dapth 3 network, where $T$ represents the length of the input time series, while $C$ denotes the number of classes for classification, as per Section III. Although this architecture appears to involve two layers between input and output, in KANs depth refers to the number of trainable connections between layers. Thus, although there are two hidden layers, the model has three learnable transformations. A depth 4 variant includes an additional hidden layer with 40 nodes. We retained the spline order $k=3$ and a fixed grid size $G=5$.
In contrast, the MLP architecture is configured as [$T$, 300, 300, 300, $C$], representing a depth 3 network. The depth 4 variant adds a hidden layer of $300$ nodes.
@@ IV-D Hyperparameter impact analysis
To provide deeper insights into the KAN architecture’s performance and provide design guidelines, we conduct an analysis of hyperparameters and configuration parameters with the goal of finding the best model for time series classification on the UCR behcmark , using a curated subset of 117 UCR datasets. To assess hyperparameter effects, we systematically varied grid size ($G$), network depth ($L$) and layer width ($M$). Table II summarizes these parameters, where the first column lists each hyperparameter and the second indicates the range of values evaluated. For each of these evaluations, the models were trained using different learning rates $lr\in\{0.0001,0.001,0.01,0.1,1\}$.
@@
In summary, to ensure robust and representative results, separate models were trained for each of the 117 UCR datasets, with input and output layer sizes matched to the the specific time series lengths and number of classes, respectively. Multiple hyperparameter settings were tested for each one, including different learning rates, depths, hidden layer sizes, and spline grid resolutions. Taken together, this resulted in the development and evaluation of several tens of thousands of individual models across the full range of datasets and configurations within the UCR archive.
Hyperparameters Values
Grid size $G\in\{3,5,10,15,20\}$
Network depth $L\in\{2,3,\ldots,10\}$
Hidden layer size $M\in\{5,10,\ldots,100\}$
Learning rate $lr\in\{0.0001,0.001,0.01,0.1,1\}$
Random seed $\{0,1,2,5,42\}$
Batch size $16$
Epochs $500$
TABLE II: Hyperparameters utilised for impact analysis for KAN and Efficient KAN models.
@@ IV-E Complexity analysis
In this section we evaluate resource consumption by calculating the average FLOPs over the 117 UCR datasets for one prediction, and the theoretical energy consumption (TEC) per prediction. FLOPs were computed theoretically using Eqs. 8 and 10 outlined in Section III. As per Section IV-A, $T\in\{16,\ldots,8,926\}$, $C\in\{2,\ldots,60\}$.
We estimated TEC as $\text{TEC}=\frac{\text{FLOPs}}{\text{FLOPS/Watt}}$ [30], where FLOPS/Watt represents the number of floating-point operations executed per second per watt. As the experiments were conducted on an NVIDIA A100 80GB PCIe GPU, its theoretical power consumption is 65 GFLOPS/Watt, for float32 operations, which were used for calculation.
@@ IV-F Interpretability analysis
Furthermore, to validate the inherent interpretability of KAN and to extract any insight from the MLP, global SHapley Additive exPlanations (SHAP) values [31] were used. SHAP offers a model-agnostic framework for global interpretability by quantifying the average contribution of each feature to the model’s predictions across the entire dataset. The contribution of a feature $x_{i}$ is represented by its Shapley value, $\phi_{i}$, which is computed based on cooperative game theory. For a model the prediction for an instance $X$ is expressed as:
$SHAP=\phi_{0}+\sum_{i=1}^{M}\phi_{i},$ (11)
where $\phi_{0}$ is the model’s baseline expected value across the dataset, $\phi_{i}$ is the Shapley value of feature $x_{i}$, and M represents the total number of features. By aggregating these values, SHAP enables insights into the global decisions of the model. In addition, SHAP is also utilized to compare the feature importance of KANs to the MLP architecture.
@@
TABLE III: Performance of reference regression architectures introduced in [23].
@@ V-A Classification performance analysis of reference architectures
Table III presents the results of the classifiers with rows representing different models and their configurations. The first two rows provide the results for MLP models and the subsequent four rows for KAN and Efficient KAN models. The first column lists the type of model, while the second column details the specific configuration of each model, including the nodes per layer and grid size for KAN models. The subsequent columns provide the performance metrics for each model, split into three main categories: precision, recall, and F1 score. Each of these categories is further divided into two subcolumns that show the mean and standard deviation (StdDev) of the results across multiple (i.e. 5) runs. The last column presents average time, measured in seconds, for training one model.
The MLP models demonstrate relatively high and consistent performance across the precision, recall, and F1 score metrics for both depths. For the 3-layer MLP, the mean precision is 0.73 with a standard deviation of 0.20, and the 4-layer model exhibits the same mean and standard deviation for precision. In terms of recall, the mean values are very similar, at 0.66 for the 3-layer model and 0.65 for the 4-layer model, with standard deviations of 0.23 and 0.20, respectively. This indicates that increasing the depth has little effect on the model’s ability to identify relevant instances. Additionally, both models maintain comparable F1 score means of 0.64 (3 layers) and 0.62 (4 layers), although the standard deviation increases slightly from 0.25 to 0.26, suggesting a minor decrease in consistency for the F1 score with the deeper model.
The original KAN implementations show a substantial drop in performance compared to the MLP models. The 3-depth KAN has a mean precision of 0.38, with a recall of 0.33 and an F1 score of 0.30. The 4-depth KAN model continues this trend with slightly lower performance, showing a mean precision of 0.37, recall of 0.32, and F1 score of 0.29. The further decline in performance suggests that adding depth does not improve generalization but may instead lead to overfitting, especially across multiple diverse datasets. These results indicate that the original KAN struggles to match the classification performance of the MLP. While previous research demonstrated that KAN performed well for time series prediction and exhibited strong generalization abilities [23], here KAN’s transferability to a different problem, i.e. classification, appears less successful.
Lower performance results come with a significantly reduced training time of $\sim 45$ seconds for depth 3 KAN, due to the GPU acceleration, as described in Section IV-E, suggesting a trade-off between performance and efficiency. However, the training time in depth 4 KAN increases to $\sim 57$ seconds, reflecting the impact of the added depth on computational cost.
@@
TABLE IV: Comparison of performance and computational characteristics across the reference models
The models trained with the Efficient KAN implementation offer a notable improvement in performance compared to the original KAN implementation. The Efficient KAN with 3-depth demonstrates a significant boost in performance with a mean precision of 0.72, recall of 0.70, and F1 score of 0.70. Similarly, the 4-depth Efficient KAN, maintains high performance metrics with a mean precision of 0.71, recall of 0.70, and F1 score of 0.69. Improvements in comparison to KAN can be attributed to implementation changes, as discussed in Section IV-C.
@@ V-B Complexity analysis for the reference models
Table IV presents the computational complexity analysis for the MLP and KAN configurations. The first two columns, listing the model and configuration, are repeated from Table III. The subsequent columns provide key complexity metrics: the third column displays the total number of lernable parameters in each model, while the fourth column reports the average theoretical FLOPs for one prediction across all used UCR datasets, based on model architectures discussed in Section III. The last column presents the Theoretical Energy Consumption (TEC) in Joules, calculated based on the FLOPs and GPU efficiency, as per Section IV-E.
In the table we observe that MLPs, with approximately $344k$ and $434k$ learnable parameters for $3$-depth and $4$-depth, respectively, are the most computationally demanding models in terms of learnable parameters and training duration. Their theoretical FLOPs reach $688,420$ and $868,720$, while requiring $1.008\times 10^{-5}$ and $1.336\times 10^{-5}$ Joules of energy per single prediction, positioning them as the best-performing models in theoretical FLOPs and TEC (Total Energy Consumption).
Next, the table reveal that the original KAN implementation has noticeably fewer learnable parameters compared to the MLP models, consistent with the theoretical analysis in Section III-A. Despite the reduction in parameters, they come with significantly higher theoretical FLOPs, with values of $6,146,993$ for depth 3 and $6,566,993$ for depth 4, due to B-spline computation, as per Eq. 8. The energy required for a single prediction also reflects this increase, amounting to $9.457\times 10^{-5}$ Joules for depth 3 and $10.103\times 10^{-5}$ Joules for depth 4. These results illustrate a trade-off, where KAN achieves a smaller parameter count but requires greater computational resources in terms of FLOPs and energy per prediction.
The last two rows for the Efficient KAN implementation show the same number of weights and theoretical FLOPs as the original KAN implementation, as the model architectures are identical. Despite higher energy consumption, Efficient KAN maintains fast training times. Along with the discussed increased performance, this balance of high performance and lower computational costs makes these models strong candidates for applications requiring both efficiency and effectiveness.